If the ratio of the roots lx^2 -nx +n= 0 is p:q
then prove  p/ q +  q/ p –  n l = 0

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
a + b = n / l
 
ab = n/l
 
a^2 + b^2 + 2ab = n^2 / l^2
 
a / b + b/a + 2 = n / l
 
p / q + q / p + 2 = n / l